OpenStudy (anonymous):

Find the first six terms of the sequence. a1 = -8, an = 5 • an-1 -8, -40, -200, -1000, -5000, -25,000 -8, -40, -35, -30, -25, -20 0, 5, -40, -35, -30, -25 -40, -200, -1000, -5000, -25,000, -125,000

OpenStudy (freckles):

-8 is the first term ...and your second term can be found by multiplying 5 to your previous term

OpenStudy (freckles):

\[a_n=5 \cdot a_{n-1} \text{ this says \to find next term just multiply previous term by 5}\]

OpenStudy (anonymous):

so the second term would be -40? because 5 x -8 = -40?

OpenStudy (freckles):

\[\text{ replace } n \text{ with 2} \\ a_2=5 \cdot a_{2-1} \\ a_2=5 \cdot a_1 \] see you are just multiplying previous term by 5 to find the term

OpenStudy (freckles):

yep

OpenStudy (anonymous):

ok thank you!

OpenStudy (freckles):

np

OpenStudy (freckles):

so what will you do for third term :)

OpenStudy (freckles):

5 times previous right?